Exercise 5-15 Operating Leverage [LO5-1, LO5-8]

Magic Realm, Inc., has developed a new fantasy board game. The company sold 27,000 games last year at a selling price of $69 per game. Fixed expenses associated with the game total $450,000 per year, and variable expenses are $49 per game. Production of the game is entrusted to a printing contractor. Variable expenses consist mostly of payments to this contractor.

Required:

1-a. Prepare a contribution format income statement for the game last year.

1-b. Compute the degree of operating leverage.

2. Management is confident that the company can sell 34,020 games next year (an increase of 7,020 games, or 26%, over last year). Given this assumption:

a. What is the expected percentage increase in net operating income for next year?

b. What is the expected amount of net operating income for next year? (Do not prepare an income statement; use the degree of operating leverage to compute your answer.)
Answer #1

1a) Income statement

Total Per unit
Sales 1863000 69
Variable cost 1323000 49
Contribution margin 540000 20
Fixed cost 450000
Net income 90000

1b) Degree of operating leverage = Contribution margin/Net income = 540000/90000 = 6

2a) Expected percentage of net income = 26*6 = 156%

2b) Amount of net operating income = 90000*2.56 = 230400